# GlitchProber: Advancing Effective Detection and Mitigation of Glitch Tokens in Large Language Models

### v1

@@ Abstract.
In this work, we aim to enhance the understanding of glitch tokens and propose techniques for their detection and mitigation. We first reveal the characteristic features induced by glitch tokens on LLMs, which are evidenced by significant deviations in the distributions of attention patterns and dynamic information from intermediate model layers. Based on the insights, we develop GlitchProber, a tool for efficient glitch token detection and mitigation. GlitchProber utilizes small-scale sampling, principal component analysis for accelerated feature extraction, and a simple classifier for efficient vocabulary screening. Taking one step further, GlitchProber rectifies abnormal model intermediate layer values to mitigate the destructive effects of glitch tokens. Evaluated on five mainstream open-source LLMs, GlitchProber demonstrates higher efficiency, precision, and recall compared to existing approaches, with an average F1 score of 0.86 and an average repair rate of 50.06%. GlitchProber unveils a novel path to address the challenges posed by glitch tokens and inspires future research toward more robust and interpretable LLMs.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
In the field of Natural Language Processing (NLP), large language models (LLMs) like GPT-4 (Achiam et al., 2023), Gemini (Team et al., 2023; Reid et al., 2024), and Claude 3 (Anthropic, 2024) have demonstrated near-human-level text generation capabilities. Their exceptional performance has led to widespread adoption (Wu et al., 2023; Topsakal and Akinci, 2023). When using these models, users provide a prompt, which the model’s tokenizer breaks down into a series of discrete tokens. These tokens are the fundamental units of information processing for the model, playing a crucial role in the usage of LLMs. Recent research (Rumbelow and mwatkins, 2023; mwatkins and Rumbelow, 2023a, b; Geiping et al., 2024; mwatkins, 2023a; Fell, 2023), however, has shown that some “glitch tokens” exist in the vocabulary of LLMs. Once included in a prompt, these special tokens can potentially lead to model errors, such as misunderstanding user intent, refusing to answer, or generating irrelevant or harmful text. Therefore, thorough analysis and detection of these glitch tokens are crucial to ensure the reliability and safety of LLMs.
@@
From these findings, we gain an insight that the glitch tokens can be efficiently detected due to the deviated distributions of intermediate layers’ outputs caused by them, and accordingly their impact can be effectively mitigated by proactively rectifying those abnormal outputs. Based on this insight, we propose a new method for glitch token detection and fix called GlitchProber. For glitch token detection, GlitchProber first samples a small subset of manually labeled glitch tokens as the sample set, and extracts the outputs of these tokens from the intermediate layers, specifically, the attention scores of the attention patterns and MLP status. It then applies Principal Component Analysis (PCA) (Abdi and Williams, 2010) dimensionality reduction to the intermediate layers’ outputs and obtains a feature representation matrix for the sample set. This matrix, along with the corresponding class labels, is used to train a Support Vector Machine (SVM) (Cortes and Vapnik, 1995) classifier, which can subsequently be employed for glitch token detection. To fix the glitch tokens, GlitchProber analyzes the activation value range of normal tokens in the intermediate MLP status and rectifies the activation states of glitch tokens. Specifically, it aims to adjust the activation patterns of glitch tokens to be closer to those of normal tokens, and thereby minimize their impact on the model’s output.
@@ 3.1. Experiment Setup
We employ a unified approach to determine whether each token is glitchy or normal in this study. While the symptoms of glitch tokens may vary across different tasks, we consistently utilize a repetition task to construct input sequences for glitch token identification. Specifically, we formulate a prompt for a repetitive task such as “Can you repeat the token ‘$\{token\}$’ and return it back to me?” This prompt is then fed into the model to assess its ability to accurately reproduce the original token. If the model fails to correctly repeat the token in its output, we classify it as a glitch token. Conversely, if the model successfully repeats the token, we categorize it as a normal token. We traverse all the 32,000 tokens of the Llama2 model and eventually identify 6,425 glitch tokens from the entire vocabulary.
To precisely capture intermediate layer outputs within the model, we resort to a transformer mechanistic interpretability tool named Transformer-lens (Nanda N, 2022). Its hook technique enables real-time access of the activation values at all layers and allows code insertion into specific intermediate layers of the model. In this study, we insert hooks into all intermediate layers during the first forward of the tested model. This approach is chosen because the first forward comprehensively reflects the model’s understanding of the input sequence and highlights the differences between normal and glitch tokens.
@@ 7.3. Glitch Token Mitigation By Fine-tuning
By constructing a dataset with questions and answers for the repetition task, we attempt to mitigate the glitch token phenomenon by fine-tuning LLMs. However, fine-tuning LLMs alters the parameters of the model, potentially compromising its basic abilities. For example, we fine-tune the Llama-2-7b-chat with a dataset containing 3,000 Q&A pairs of repetition tasks. To evaluate the model’s basic skills, we use three widely accepted datasets namely GSM8K (Cobbe et al., 2021), HumanEval (Chen et al., 2021), and MMLU (Hendrycks et al., 2020). Detailed results presented on our website (GlitchProber, 2024) indicate that the model’s ability in code writing and solving math problems post fine-tuning is significantly diminished compared to the original model. In contrast, GlitchProber modifies the process of model calculation without altering the model parameters, thereby preserving the basic abilities of LLMs.

### v2

@@ Abstract.
In this work, we aim to enhance the understanding of glitch tokens and propose techniques for their detection and mitigation. We first reveal the characteristic features induced by glitch tokens on LLMs, which are evidenced by significant deviations in the distributions of attention patterns and dynamic information from intermediate model layers. Based on the insights, we develop GlitchProber, a tool for efficient glitch token detection and mitigation. GlitchProber utilizes small-scale sampling, principal component analysis for accelerated feature extraction, and a simple classifier for efficient vocabulary screening. Taking one step further, GlitchProber rectifies abnormal model intermediate layer values to mitigate the destructive effects of glitch tokens. Evaluated on five mainstream open-source LLMs, GlitchProber demonstrates higher efficiency, precision, and recall compared to existing approaches, with an average F1 score of 0.86 and an average repair rate of 50.06%. GlitchProber unveils a novel path to address the challenges posed by glitch tokens and inspires future research toward more robust and interpretable LLMs. Our code is available at https://github.com/LLM-Integrity-Guard/GlitchProber.
LLM security, Glitch token, LLM analysis
journalyear: 2024copyright: acmlicensedconference: 39th IEEE/ACM International Conference on Automated Software Engineering; October 27-November 1, 2024; Sacramento, CA, USAbooktitle: 39th IEEE/ACM International Conference on Automated Software Engineering (ASE ’24), October 27-November 1, 2024, Sacramento, CA, USAdoi: 10.1145/3691620.3695060isbn: 979-8-4007-1248-7/24/10ccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
In the field of Natural Language Processing (NLP), large language models (LLMs) like GPT-4 (Achiam et al., 2023), Gemini (Team et al., 2023; Reid et al., 2024), and Claude 3 (Anthropic, 2024) have demonstrated near-human-level text generation capabilities. Their exceptional performance has led to widespread adoption (Wu et al., 2023; Topsakal and Akinci, 2023; Guan et al., 2024). When using these models, users provide a prompt, which the model’s tokenizer breaks down into a series of discrete tokens. These tokens are the fundamental units of information processing for the model, playing a crucial role in the usage of LLMs. Recent research (Rumbelow and mwatkins, 2023; mwatkins and Rumbelow, 2023a, b; Geiping et al., 2024; mwatkins, 2023a; Fell, 2023), however, has shown that some “glitch tokens” exist in the vocabulary of LLMs. Once included in a prompt, these special tokens can potentially lead to model errors, such as misunderstanding user intent, refusing to answer, or generating irrelevant or harmful text. Therefore, thorough analysis and detection of these glitch tokens are crucial to ensure the reliability and safety of LLMs.
@@
From these findings, we gain an insight that the glitch tokens can be efficiently detected due to the deviated distributions of intermediate layers’ outputs caused by them, and accordingly their impact can be effectively mitigated by proactively rectifying those abnormal outputs. Based on this insight, we propose a new method for glitch token detection and fix called GlitchProber. For glitch token detection, GlitchProber first samples a small subset of manually labeled glitch tokens as the sample set, and extracts the outputs of these tokens from the intermediate layers, specifically, the attention scores of the attention patterns and MLP status. It then applies Principal Component Analysis (PCA) (Abdi and Williams, 2010) dimensionality reduction to the intermediate layers’ outputs and obtains a feature representation matrix for the sample set. This matrix, along with the corresponding class labels, is used to train a Support Vector Machine (SVM) (Cortes and Vapnik, 1995) classifier, which can subsequently be employed for glitch token detection. To fix the glitch tokens, GlitchProber analyzes the activation value range of normal tokens in the intermediate MLP status and rectifies the activation states of glitch tokens. Specifically, it aims to adjust the activation patterns of glitch tokens to be closer to those of normal tokens, and thereby minimize their impact on the model’s output. Our work is published on our website (GlitchProber, 2024).
@@ 3.1. Experiment Setup
We employ a unified approach to determine whether each token is glitchy or normal in this study based on existed definition of glitch tokens (Fell, 2023). While the symptoms of glitch tokens may vary across different tasks, we consistently utilize a repetition task to construct input sequences for glitch token identification.
In the context of our work, a repetitive task refers to a specifically designed experimental procedure to test the fidelity of a language model in reproducing input tokens. This task is utilized primarily for the identification and categorization of tokens based on their performance when repetitively prompted. Specifically, the repetitive task involves the following steps:
  1. (1)
    Formulating a prompt that requires the model to duplicate a specific token. The typical prompt structure is, “Can you repeat the token ‘$\{token\}$’ and return it back to me?” This format is deliberately chosen to minimize contextual influence and focus purely on the token reproduction capability of the model.
  2. (2)
    Submitting the prompt to the model, which is configured by setting the temperature parameter to zero.
  3. (3)
    Observing and analyzing the model’s output to determine whether it accurately reproduces the input token. The output is deemed successful if the model returns the exact token as requested; otherwise, the token is classified as a glitch token.
We traverse all the 32,000 tokens of the Llama2 model and eventually identify 6,425 glitch tokens from the entire vocabulary.To precisely capture intermediate layer outputs within the model, we resort to a transformer mechanistic interpretability tool named Transformer-lens (Nanda N, 2022). Its hook technique enables real-time access of the activation values at all layers and allows code insertion into specific intermediate layers of the model. In this study, we insert hooks into all intermediate layers during the first forward of the tested model. This approach is chosen because the first forward comprehensively reflects the model’s understanding of the input sequence and highlights the differences between normal and glitch tokens.
@@ 7.1. Hyperparameters Choice of GlitchProber
We further examine the rationale behind the selection of the hyperparameter $\gamma$, which determines the sampling rate from the model’s token vocabulary $V$ for the sample set $S$ in GlitchProber. The parameter $\gamma$ is pivotal in balancing detection accuracy against computational efficiency. Empirical analysis, as shown in Table 7, demonstrates that a $\gamma$ increment from 0.1 to 0.3 improves recall from 0.6922 to 0.7457 and raises the F1 score from 0.8181 to 0.8543, with a reasonable increase in computational time. However, further increasing $\gamma$ to 0.7, while boosting recall and F1 scores to 0.7812 and 0.8772 respectively, significantly extends processing times to over 100 minutes. Therefore, a $\gamma$ range of 0.1 to 0.3 is recommended, as it optimally balances performance gains with computational efficiency, ensuring that GlitchProber remains practical for operational use.
Table 7. Model performance across different gamma values on Llama2 model
Sampling Rate $\gamma=0.1$ $\gamma=0.2$ $\gamma=0.3$ $\gamma=0.5$ $\gamma=0.7$
Precision 100% 100% 100% 100% 100%
Recall 69.22% 72.81% 74.57% 76.65% 78.12%
F1 Score 81.81% 84.26% 85.43% 86.78% 87.72%
Time 61min 38s 68min 14s 74min 30s 86min 52s 100min 41s
@@ 7.3. Glitch Token Mitigation By Fine-tuning
Table 8. Performance comparison of original and modified model using Finetuning and GlitchProber for Mitigation
Dataset Original Model GlitchProber Finetuning
GSM8K 0.315 0.301 0.238
HumanEval pass@1 0.129 0.103 0.009
HumanEval pass@5 0.190 0.161 0.024
MMLU 0.453 0.417 0.229
GlitchProber adaptively modifies the process of model calculation without altering the model parameters, thereby minimizing the mitigation impact on model performance and preserving the basic abilities of LLMs. In contrast, we also construct a dataset with Q&A for the repetition task and attempt to mitigate the glitch token phenomenon by fine-tuning LLMs. However, compared with GlitchProber, fine-tuning LLMs alters the parameters of the model, potentially compromising its basic abilities. For example, we fine-tune the Llama-2-7b-chat with a dataset containing 3,000 Q&A pairs of repetition tasks. To evaluate the model’s basic skills, we use three widely accepted datasets namely GSM8K (Cobbe et al., 2021), HumanEval (Chen et al., 2021), and MMLU (Hendrycks et al., 2020).
Detailed results presented in Table 8 indicate that the model’s ability in code writing and solving math problems post GlitchProber is comparable with the original model. and significantly better compared to the fine-tuned model, which notably diminished the basic abilities of original model.
@@ 7.4. Threats to Validity
Acknowledgement
This work was supported by the National NSF of China (grants No.62302176, No.62072046, 62302181), the Key R&D Program of Hubei Province (2023BAB017, 2023BAB079), and the Knowledge Innovation Program of Wuhan-Basic Research (2022010801010083).